# ASVD: Activation-aware Singular Value Decomposition for Compressing Large Language Models

### v3

@@ Abstract
In this paper, we introduce a new post-training compression paradigm for Large Language Models (LLMs) to facilitate their wider adoption. We delve into LLM weight low-rank factorization, and find that the challenges of this task stem from the outlier phenomenon in the LLM activations and the sensitivity difference among various kinds of layers. To address these issues, we propose a training-free approach called Activation-aware Singular Value Decomposition (ASVD). Specifically, ASVD manages activation outliers by scaling the weight matrix based on the activation distribution, thereby enhancing decomposition accuracy. Additionally, we propose an efficient iterative calibration process to optimize layer-specific decomposition by addressing the varying sensitivity of different LLM layers. ASVD can compress a network by 10-20%, without compromising the performance of LLMs. Based on the success of the low-rank decomposition of projection matrices in the self-attention module, we further introduce ASVD to compress the KV cache. By reducing the channel dimension of KV activations, memory requirements for KV cache can be largely reduced. Thanks to the 50-75% reduction in the rank of the KV projection matrices, ASVD can further achieve 50% KV cache reductions without performance drop in a training-free manner. Code and compressed models are available at ASVD4LLM.
@@ 1 Introduction
Numerous strategies have been proposed to address the memory consumption challenges in Large Language Models (LLMs), as reviewed by Zhu et al. [2023], Ding et al. [2023]. In the realm of LLM compression, various techniques have been extensively explored, including weight quantization [Dettmers et al., 2022], network pruning [Frantar & Alistarh, 2023], and knowledge distillation [Agarwal et al., 2023]. Distinct from these approaches, the paradigm of low-rank matrix decomposition is less explored but holds significant promise. Decomposition involves approximating the weight matrices in LLMs with matrices of lower rank, effectively reducing the model size. Given the massive number of parameters in LLMs, low-rank decomposition offers significant potential for memory reduction. Furthermore, low-rank decomposition can complement existing LLM compression techniques by further compressing quantized or pruned models, enhancing overall efficiency [Cheng et al., 2017]. Despite its potential, it remains a relatively underutilized approach for LLM compression.
@@
From the perspective of network compression, traditional low-rank decomposition methods typically adhere to a straightforward process: initially training the original model and subsequently fine-tuning the decomposed model [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021, Hsu et al., 2022]. While this approach is effective, it is resource-intensive and requires the entire training dataset and substantial computational power for end-to-end backpropagation. Applying this method to LLMs would encounter major challenges. Firstly, the training data for LLMs may not always be readily available, often restricted by privacy and commercial considerations. Secondly, the training process for these models is notoriously expensive, both in terms of time and computational resources. Given these constraints, the concept of “training-free” compression emerges as a more viable approach for LLMs [Zhu et al., 2023]. This approach includes methods like LLM post-training quantization [Dettmers et al., 2022, Yuan et al., 2023] and LLM post-training pruning [Frantar & Alistarh, 2023], which compress LLMs without the need for extensive retraining. These training-free (i.e., post-training) methods offer a more practical solution for efficiently compressing LLMs.
@@
To realize LLM low-rank decomposition in a training-free manner, we conduct an extensive analysis of the baseline methods for LLM decomposition. We first observe that straightforward application of existing low-rank decomposition techniques, which typically necessitate training, turns out ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021]. Digging into the failures, we reveal several challenges unique to post-training decomposition for LLMs. One key challenge involves managing outliers in the activations, which significantly intensifies the decomposition error. This challenge echoes findings in LLM quantization research [Lin et al., 2023, Kim et al., 2023], highlighting the importance of handling such outliers in LLMs.
@@
To address these challenges, we propose Activation-aware Singular Value Decomposition (ASVD), where the distribution of activations are considered into the weight decomposition process. Specifically, it achieves this by scaling the values in the weight matrix column-wisely, based on the distribution patterns observed across input and output channels. This adjustment proves particularly beneficial for channels with outlier activations, allowing the ASVD to allocate enhanced focus to these specific weights. Such targeted adjustment facilitates a more accurate reconstruction of the weight matrix, thereby minimizing prediction errors and improving the overall efficiency of the decomposition process. We further investigate the varying sensitivity of different LLM layers to decomposition. We find that weights in Multi-Head Attention layers  [Vaswani et al., 2017] tend to be more resilient to decomposition compared to those in Multi-Layer Perceptron layers. This sensitivity variability across layers prompts us to develop a method to assign the compression ratio for each layer. ASVD assesses each layer’s sensitivity to decomposition at different ranks, enabling us to assign a suitable rank for optimal decomposition. Note that this probing assess is very efficient, requiring only a limited sample set for evaluation.
Empirically, we evaluate ASVD with LLaMA and LLaMA-2 [Touvron et al., 2023a, b]. Our experiments reveal that ASVD can reduce the rank of the weight matrix by 50-75% (corresponds to 10-20% compression ratio), with only a marginal 1% accuracy loss on benchmarks like MMLU [Hendrycks et al., 2020]. We also validate ASVD is compatible with 4/6/8-bit quantization, as shown in Fig. 1(a).
Importantly, leveraging the successful low-rank decomposition of projection matrices in the self-attention module, we can integrate ASVD with KV cache compression. Specifically, by applying ASVD to decompose the Key/Value projection matrices, we can derive low-rank intermediate activations that serve as replacements for the KV cache stored in high-dimension space as shown in Fig. 1(b). This substitution significantly reduces the memory usage of the KV cache, enabling support for larger batch sizes or longer sequence lengths, which are essential for real-world applications [Yuan et al., 2024]. In practice, by replacing the KV cache with low-rank intermediate activations, we can reduce up to 50% of the memory consumption of KV cache.
@@ 2 Related Work
Large Language Model Compression. The field of model compression for Large Language Models (LLMs) has seen a surge of innovative techniques aimed at mitigating the substantial computation and memory requirements these models demand [Zhu et al., 2023, Yuan et al., 2024]. Various methods have emerged to address this challenge, each taking a unique approach to reduce the memory footprint of LLMs. These methods primarily fall into three categories: weight quantization [Courbariaux et al., 2015, Dettmers et al., 2022], network pruning [LeCun et al., 1989, Frantar & Alistarh, 2023], and knowledge distillation [Hinton et al., 2015, Agarwal et al., 2023]. For the wide body of research on LLM compression, please refer to [Zhu et al., 2023] for the comprehensive survey. Among these methods, weight quantization has gained significant traction in the context of LLMs due to its effectiveness. However, despite its popularity as a neural network compression technique, low-rank factorization has not been extensively explored in the realm of LLMs.
@@
Low-rank Decomposition. In the realm of low-rank decomposition [Schotthöfer et al., 2022] for neural network compression, existing methods can be broadly classified into two categories: fixed low rank and variable low rank approaches. Fixed rank methods typically involve decomposing weight matrices of pre-trained networks using techniques like Singular Value Decomposition (SVD) or tensor decomposition, followed by fine-tuning the factorized network [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015]. They also involve constraining weight matrices to maintain a fixed low rank during training [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021], or constructing layers as linear combinations of layers with varying ranks [Ioannou et al., 2015]. A notable limitation of these methods is the introduction of matrix decomposition rank as an additional hyperparameter requiring fine-tuning. In contrast, rank-adaptive methods address this limitation by automatically determining and adjusting the low-rank structure. In particular, Kim et al. [2015, 2019] apply heuristics search to pre-determine the decomposition rank, while Wen et al. [2017] learn low-rank weights through a loss function penalizing approximated matrix ranks. Li et al. [2023] use low-rank approximation plus a sparse matrix to compress the weight matrix in transformers.
@@
However, none of these methods have worked in the era of LLMs due to their training-require nature. In this work, we introduce ASVD, a post-training LLM decomposition approach enabling the adaptive determination of SVD ranks to optimize the matrix approximations based on feature activations. To our knowledge, this is the first attempt to compress the weights of LLMs through decomposition in a training-free manner.
@@ 3.1 Naïve SVD for Compressing Weight Matrix
Singular Value Decomposition (SVD) can be used to decomposing the weights of linear layers, which involves decomposing a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into three matrices: $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}^{T}$, such that $\mathbf{W}\approx\mathbf{W}\mathbf{\Sigma}\mathbf{V}^{T}$), where $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix, the diagonal values in $\mathbf{\Sigma}$ are the singular values of $\mathbf{W}$, and $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are corresponding right and left singular vector matrices, respectively [Demmel, 1997].
The SVD compression process for a weight matrix can be summarized in three steps: Decomposition: Factorize $\mathbf{W}$ using SVD. Truncation: Retain the top $k$ singular values and their corresponding right and left singular vectors. This results in approximated matrices $\mathbf{U}_{k}$, $\mathbf{\Sigma}_{k}$, and $\mathbf{V}_{k}^{T}$, where the right singular vector matrix $\mathbf{U}_{k}$ is $m\times k$, singular $\mathbf{\Sigma}_{k}$ is $k\times k$, and left singular vector matrix $\mathbf{V}_{k}^{T}$ is $k\times n$. The choice of $k$ is critical in balancing the compression ratio and the compressed model’s performance. Reconstruction: Reconstruct an approximated weight matrix: $\mathbf{W}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}$.
@@ 3.2 Challenges of Compressing LLMs via SVD
$\mathbf{W}_{k}^{\star}$ $=\arg\min_{\mathbf{W}_{k}}\mathcal{L}(\mathbf{W}_{k}),$ (1)
$\arg\min_{\mathbf{W}_{k}}$ $\mathcal{L}(\mathbf{W}_{k})=\|\mathbf{W}_{k}\mathbf{X}-\mathbf{W}\mathbf{X}\|$
@@
Challenge 2: Singular Values Variations among Layers: The distribution of singular values within a matrix is indicative of its sparsity and, by extension, its sensitivity to certain types of information [Kim et al., 2015, 2019, Wen et al., 2017]. In LLMs, there is a notable variation in singular values across different layers. Specifically, some layers exhibit a concentration of large singular values, signifying less sensitivity to weight variation. This characteristic often correlates with these layers being easy to compress. Conversely, other layers in the LLMs display a more uniform distribution of smaller singular values. Such a pattern suggests a balanced contribution from various singular vector pairs. This variability in the distribution of singular values among layers presents a unique challenge, as it implies that each layer may require a tailored approach to decompose and maintain the overall functionality of the LLM.
@@ 3.3 ASVD: Activation-aware Singular Value Decomposition
Our ASVD is designed to refine the weight matrix $\mathbf{W}$ in LLMs by taking into account the effect of input activation channels. The process comprises the following three steps:
Scaling the Weight Matrix. The first step involves scaling the weight matrix $\mathbf{W}$ by a diagonal matrix $\mathbf{S}$. This matrix $\mathbf{S}$ is constructed to represent the impact of input channels on the weights, essentially adjusting $\mathbf{W}$ to better adapt with the activation patterns of the input $\mathbf{X}$, as demonstrated in Fig. 2 (Bottom line). The scaled weight matrix is denoted as $\mathbf{WS}$, and the equation is formulated as:
@@
This scaling step is crucial as it modifies the weight matrix to reflect the varying significance of different input channels, making the decomposition more effective and targeted.
Applying SVD to the Scaled Matrix. After scaling the weight matrix, the next step is to apply SVD to the scaled matrix $\mathbf{WS}$. The SVD of $\mathbf{WS}$ is expressed as $\mathbf{WS}=\mathbf{U}^{\prime}\mathbf{\Sigma}^{\prime}{\mathbf{V}^{\prime}}^{T}$. To realize the matrix compression, we truncate this decomposition to retain only the first $k$ singular values and corresponding vectors. The truncated form of the decomposition is represented as:
@@
This step is pivotal as it ensures that the most significant aspects of the scaled weight matrix are retained. While less critical information, which contributes minimally to the model’s output, is discarded.
Reconstructing the Approximated Weight Matrix. The final step in the ASVD process is to reconstruct an approximation of the original weight matrix. This reconstruction is accomplished by multiplying the truncated decomposition $\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime}}^{T}$ with the inverse of the scaling matrix $\mathbf{S}^{-1}$. The multiplication of ${\mathbf{V}_{k}^{\prime}}^{T}$ with $\mathbf{S}^{-1}$ yields a new matrix ${\mathbf{V}_{k}^{\prime\prime}}^{T}$ (i.e., ${\mathbf{V}_{k}^{\prime\prime}}^{T}={\mathbf{V}_{k}^{\prime}}^{T}\mathbf{S}^{-1}$), which is the transformed version of the truncated right singular vectors considering the scaling effect. As a result, the final approximated weight matrix can be expressed as:
$\mathbf{W}\approx\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime\prime}}^{T}.$ (5)
ASVD focus on channel-wise activation outliers, striving to mitigate these outliers by appropriately scaling the corresponding weights. In this way, the activation variation can be mitigated as illustrated in Fig. 2. Note that the diagonal nature of $\mathbf{S}$ simplifies its inversion process, making $\mathbf{S}^{-1}$ computationally efficient. The subsequent challenge is to devise a method or criteria for accurately determining $\mathbf{S}$ so that it effectively aligns with the activation patterns observed in the LLM, ensuring that the scaling process optimally addresses the outlier activations.
$S_{ii}$ is pivotal in moduling the input channel as: $(WS)_{:,i}=W_{:,i}S_{ii}$
Setting Scaling Matrix $\mathbf{S}$. Each diagonal element in the scaling matrix $\mathbf{S}$, particularly $\mathbf{S}_{ii}$ is pivotal in moduling the input channel as: $(\mathbf{WS})_{:,i}=\mathbf{W}_{:,i}\mathbf{S}_{ii}$. Each entry of $\mathbf{S}_{ii}$ specifically scales the influence of the $i$-th input channel. This scaling is instrumental in adjusting how each activation channel impacts the weight matrix during the decomposition process. The significance of $\mathbf{S}_{ii}$ is visualized in Fig.2 (bottom line), where its role in modulating channel influence is depicted. We use absolute mean value of input activation to set $\mathbf{S}_{ii}$. We calculates $\mathbf{S}_{ii}$ by calculating the absolute mean value of the activations in the $i$-th channel. Mathematically, if $\mathbf{X}_{i,:}$ represents the input activation values of the $i$-th channel, then $\mathbf{S}_{ii}$ is computed as: $\mathbf{S}_{ii}=(\frac{1}{n}\sum_{j=1}^{n}|\mathbf{X}_{ij}|)^{\alpha}$, where $n$ is the total number of activations for the $i$-th channel and hyper-parameter $\alpha$ provides flexibility to adjust the level of activation sensitivity incorporated into the scaling. This method focuses on the average magnitude of activation in each channel, capturing the general intensity of activation signals irrespective of their positive or negative nature.
3.4 KV Cache Compression

### v4

@@ Abstract
In this paper, we introduce a new post-training compression paradigm for Large Language Models (LLMs) to facilitate their wider adoption. We delve into LLM weight low-rank decomposition, and find that the challenges of this task stem from the distribution variance in the LLM activations and the sensitivity difference among various kinds of layers. To address these issues, we propose a training-free approach called Activation-aware Singular Value Decomposition (ASVD). Specifically, ASVD manages activation outliers by transforming the weight matrix based on the activation distribution. This transformation allows the outliers in the activation matrix to be absorbed into the transformed weight matrix, thereby enhancing decomposition accuracy. Additionally, we propose an efficient iterative calibration process to optimize layer-specific decomposition by addressing the varying sensitivity of different LLM layers. In this way, ASVD can compress a network by 10%-30%. Based on the success of the low-rank decomposition of projection matrices in the self-attention module, we further introduce ASVD to compress the KV cache. By reducing the channel dimension of KV activations, memory requirements for KV cache can be largely reduced. ASVD can further achieve 50% KV cache reductions without performance drop in a training-free manner. Code is anonymously available in supplementary materials.
@@ 1 Introduction
In the realm of Large Language Models (LLMs) compression, various techniques have been extensively explored, including weight quantization [Dettmers et al., 2022], network pruning [Frantar & Alistarh, 2023], and knowledge distillation [Agarwal et al., 2023]. Distinct from these approaches, the paradigm of low-rank matrix decomposition is less explored in LLMs but holds significant promise. Decomposition involves approximating the weight matrices in neural networks with matrices of lower rank, effectively reducing the model size. Given the massive number of parameters in LLMs, low-rank decomposition offers significant potential for memory reduction. Furthermore, low-rank decomposition can complement existing LLM compression techniques by further compressing quantized or pruned models, enhancing overall efficiency [Cheng et al., 2017].
@@
From the perspective of network compression, traditional low-rank decomposition methods typically adhere to a straightforward process: initially training the original model and subsequently fine-tuning the decomposed model [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021, Hsu et al., 2022]. While this approach is effective, it is resource-intensive and requires the entire training dataset and substantial computational power for end-to-end backpropagation. Applying this method to LLMs would encounter major challenges. Firstly, the training data for LLMs may not always be readily available, often restricted by privacy and commercial considerations. Secondly, the training process for these models is notoriously expensive, both in terms of time and computational resources.
Given these constraints, the concept of “training-free” compression emerges as a more viable approach for LLMs [Zhu et al., 2023]. This approach includes methods like LLM post-training quantization [Dettmers et al., 2022, Yuan et al., 2023] and LLM post-training pruning [Frantar & Alistarh, 2023], which compress LLMs without the need for extensive retraining. These training-free (i.e., post-training) methods offer a more practical solution for efficiently compressing LLMs.
@@
To realize LLM low-rank decomposition in a training-free manner, we conduct an extensive analysis of the baseline methods for LLM decomposition. We first observe that straightforward application of existing low-rank decomposition techniques, which typically necessitate training, turns out ineffective for LLMs [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015, Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021].
Digging into the failures, we reveal two challenges to post-training decomposition for LLMs. Managing activation distribution in LLMs: This challenge involves addressing outliers in the activations, which can intensify the decomposition error. The importance of handling such outliers in LLMs echoes findings in recent quantization research [Lin et al., 2023, Kim et al., 2023]. These outliers can disproportionately affect the accuracy of matrix approximations, leading to suboptimal compression results. Balancing layer’s decomposition sensitivity: Some layers are more sensitive to the decompostion than others, and decomposing them uniformly can lead to significant performance degradation. The key challenge is to balance the sensitivity of each layer with the efficiency of the whole network’s decomposition.
@@
Targeting challenge , we propose the activation-aware decomposition method, where the distribution of activations are considered into the weight decomposition process. Specifically, we transform the values in the weight matrix column-wisely via a scaling matrix. The scaling matrix is designed based on the distribution patterns observed across input activation channels. This adjustment proves particularly beneficial for activation with outliers, allowing the decomposition to allocate enhanced focus to these specific weights. Targeting challenge , we further investigate the varying sensitivity of different LLM layers to decomposition. We find that weights in Multi-Head Attention layers  [Vaswani et al., 2017] tend to be more resilient to decomposition compared to those in Multi-Layer Perceptron layers. This sensitivity variability across layers prompts us to develop a method to assign the compression ratio for each layer. ASVD assesses each layer’s sensitivity to decomposition at different ranks, enabling us to assign a suitable rank for optimal decomposition. Note that this probing assess is very efficient, requiring only a limited sample set for evaluation.
Our experiments reveal that ASVD can reduce the rank of the weight matrix by 10% to 90% in different layers, and it can achieve compression of model size 10%-30% in LLaMA models [Touvron et al., 2023a; b]. We also validate ASVD is compatible with 4/8-bit weight quantization, which is described in Sect. 4.4.
Importantly, leveraging the successful low-rank decomposition of projection matrices in the self-attention module, we can integrate ASVD with KV cache compression. Specifically, by applying ASVD to decompose the Key/Value projection matrices, we can derive low-rank intermediate activations that serve as replacements for the KV cache stored in a high-dimension space, as shown in Fig. 1(b). This substitution significantly reduces the memory usage of the KV cache, enabling support for larger batch sizes or longer sequence lengths, which are essential for real-world applications [Yuan et al., 2024]. In practice, by replacing the KV cache with intermediate low-rank activations, we can reduce up to 50% of the memory consumption of the KV cache.
@@ 2 Related Work
Large Language Model Compression. The field of model compression for Large Language Models (LLMs) has seen a surge of innovative techniques aimed at mitigating the substantial computation and memory requirements these models demand [Zhu et al., 2023, Yuan et al., 2024]. Various methods have emerged to address this challenge, each taking a unique approach to reduce the memory footprint of LLMs. These methods primarily fall into three categories: weight quantization [Courbariaux et al., 2015, Dettmers et al., 2022], network pruning [LeCun et al., 1989, Frantar & Alistarh, 2023], and knowledge distillation [Hinton et al., 2015, Agarwal et al., 2023]. For the wide body of research on LLM compression, please refer to [Zhu et al., 2023] for the comprehensive survey. Among these methods, weight quantization has gained significant traction in the context of LLMs due to its effectiveness. However, despite its popularity as a neural network compression technique, low-rank factorization has not been extensively explored in the realm of LLMs. Recognizing this gap, we introduce a novel low-rank decomposition method tailored specifically for decomposing the weight matrices of LLMs in a training-free manner.
@@
Low-rank Decomposition. In the realm of low-rank decomposition [Schotthöfer et al., 2022] for neural network compression, existing methods can be broadly classified into two categories: fixed low rank and variable low rank approaches. Fixed rank methods typically involve decomposing weight matrices of pre-trained networks using techniques like Singular Value Decomposition (SVD) or tensor decomposition, followed by fine-tuning the factorized network [Denton et al., 2014, Lebedev et al., 2014, Sainath et al., 2013, Moczulski et al., 2015]. They also involve constraining weight matrices to maintain a fixed low rank during training [Jaderberg et al., 2014, Khodak et al., 2021, Wang et al., 2021], or constructing layers as linear combinations of layers with varying ranks [Ioannou et al., 2015]. A notable limitation of these methods is the introduction of matrix decomposition rank as a hyperparameter requiring fine-tuning. In contrast, rank-adaptive methods address this limitation by automatically determining and adjusting the low-rank structure. In particular, Kim et al. [2015; 2019] apply heuristics search to pre-determine the decomposition rank, while Wen et al. [2017] learn low-rank weights through a loss function penalizing approximated matrix ranks. Li et al. [2023] use low-rank approximation plus a sparse matrix to compress the weight matrix in transformers.
@@
However, none of these methods have worked in the era of LLMs due to their training-require nature. We propose ASVD, a post-training LLM decomposition approach enabling the adaptive determination of SVD ranks to optimize the matrix approximations based on feature activations. To our knowledge, ASVD represents the first attempt to compress the weights of LLMs through decomposition in a training-free manner. Since the introduction of ASVD, there have been subsequent works on training-free LLM decomposition, such as SVD-LLM [Wang et al., 2024] and Palu [Chang et al., 2024]. These follow-up studies underscore the significance and potential of our approach. We hope that our proposed post-training LLM decomposition method can establish a new paradigm for LLM compression, opening up avenues for more efficient and accessible deployment of LLMs.
@@ 3.1 Naïve SVD for Compressing Weight Matrix
Singular Value Decomposition (SVD) can be used to decompose the weights of linear layers, which involves decomposing a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into three matrices: $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}^{T}$, such that $\mathbf{W}\approx\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$), where $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix, the diagonal values in $\mathbf{\Sigma}$ are the singular values of $\mathbf{W}$, and $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are corresponding right and left singular vector matrices, respectively [Demmel, 1997].
The SVD compression process for a weight matrix can be summarized in three steps: Decomposition: Factorize $\mathbf{W}$ using SVD. Truncation: Retain the top $k$ singular values and their corresponding right and left singular vectors. This results in approximated matrices $\mathbf{U}_{k}$, $\mathbf{\Sigma}_{k}$, and $\mathbf{V}_{k}^{T}$, where the right singular vector matrix $\mathbf{U}_{k}$ is $m\times k$, singular $\mathbf{\Sigma}_{k}$ is $k\times k$, and left singular vector matrix $\mathbf{V}_{k}^{T}$ is $k\times n$. The choice of $k$ is critical in balancing the compression ratio and the compressed model’s performance. Reconstruction: Reconstruct an approximated weight matrix: $\mathbf{W}_{k}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{V}_{k}^{T}$.
@@ 3.2 Challenges of Compressing LLMs via SVD
$\mathbf{W}_{k}^{\star}$ $=\arg\min_{\mathbf{W}_{k}}\|\mathbf{W}_{k}\mathbf{X}-\mathbf{W}\mathbf{X}\|_{F}^{2}.$ (1)
@@
Challenge 2: Singular Values Variations among Layers: The distribution of singular values within a matrix is indicative of its sparsity and, by extension, its sensitivity to certain types of information [Kim et al., 2015; 2019, Wen et al., 2017]. In LLMs, there is a notable variation in singular values across different layers. Specifically, some layers exhibit a concentration of large singular values, signifying less sensitivity to weight variation. This characteristic often correlates with these layers being easy to compress. Conversely, other layers in the LLMs display a more uniform distribution of smaller singular values. Such a pattern suggests a balanced contribution from various singular vector pairs. This variability in the distribution of singular values among layers presents a unique challenge, as it implies that each layer may require a tailored approach to decompose and maintain the overall functionality of the LLM.
@@ 3.3 ASVD: Activation-aware Singular Value Decomposition
ASVD is designed to refine the weight matrix $\mathbf{W}$ in LLMs by taking into account the effect of input activation channels. The process comprises the following three steps:
Transforming the Weight Matrix. The first step involves transforming the weight matrix $\mathbf{W}$ into an invertible matrix $\mathbf{S}$. The transform is denoted as $\mathbf{WS}$. Because the matrix $\mathbf{S}$ is invertible, we can have this equation:
@@
Applying SVD to the Transformed Matrix. After transforming the weight matrix, the next step is to apply SVD to the transformed matrix $\mathbf{WS}$. The SVD of $\mathbf{WS}$ is expressed as $\mathbf{WS}=\mathbf{U}^{\prime}\mathbf{\Sigma}^{\prime}{\mathbf{V}^{\prime}}^{T}$. To reduce the elements in these matrices, we truncate them to retain only the top-$k$ singular values. The truncated form of the decomposition is represented as:
@@
This step ensures that the most significant aspects of the scaled weight matrix are retained. While less critical information, which contributes minimally to the model’s output, is discarded.
Reconstructing the Approximated Weight Matrix. The final step is to reconstruct an approximation of the original weight matrix. We multiply ${\mathbf{V}_{k}^{\prime}}^{T}$ with $\mathbf{S}^{-1}$ to produce a new matrix ${\mathbf{V}_{k}^{\prime\prime}}^{T}$:
${\mathbf{V}_{k}^{\prime\prime}}^{T}={\mathbf{V}_{k}^{\prime}}^{T}\mathbf{S}^{-1}.$ (5)
Note that the matrix ${\mathbf{V}_{k}^{\prime\prime}}^{T}$ has the same shape as the matrix ${\mathbf{V}_{k}^{\prime}}^{T}$. In this way, the weight matrix can be approximated by:
$\mathbf{W}=(\mathbf{WS})\mathbf{S}^{-1}\approx(\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime}}^{T})\mathbf{S}^{-1}=\mathbf{U}_{k}^{\prime}\mathbf{\Sigma}_{k}^{\prime}{\mathbf{V}_{k}^{\prime\prime}}^{T}=\mathbf{W}_{k}.$ (6)
Setting the Transform Matrix $\mathbf{S}$. The transform matrix $\mathbf{S}$ is constructed to adjust $\mathbf{W}$ to better adapt with the activation patterns of the input $\mathbf{X}$. A simple method is to set the transform matrix as a diagonal matrix. The computation of the linear layer can be transformed by:
$\mathbf{W}\mathbf{X}=(\mathbf{WS})(\mathbf{S}^{-1}\mathbf{X}).$ (7)
Each diagonal element in the matrix $\mathbf{S}_{ii}$ transforms the $i$-th input channel of weight as: $(\mathbf{WS})_{:,i}=\mathbf{W}_{:,i}\mathbf{S}_{ii}$. Because $\mathbf{S}^{-1}$ is also a diagonal matrix, the $\mathbf{S}_{ii}^{-1}$ scales the $i$-th channel of the activation as $\mathbf{S}^{-1}_{ii}\mathbf{X}_{i,:}$. This scaling adjusts how each activation channel impacts the weight matrix during the decomposition process. We visualize the impact of the adjustment in Fig.2. We use a small number of corpus sent to the LLM and calculate the absolute mean value of input activation channel. Then we set $\mathbf{S}_{ii}$ according to the absolute mean value of the activations in the $i$-th channel:
$\mathbf{S}_{ii}:=(\frac{1}{n}\sum_{j=1}^{n}|\mathbf{X}_{ij}|)^{\alpha},$ (8)
where $n$ is the total number of activations for the $i$-th channel and hyper-parameter $\alpha$ provides flexibility to adjust the level of activation sensitivity incorporated into the scaling. This method focuses on the average magnitude of activation in each channel, capturing the general intensity of activation signals regardless of their positive or negative nature. Since we only need to do the LLM inference several times, this method is very fast.
Another method to set the transform matrix $\mathbf{S}$ to is to optimize the output error introduced by decomposition directly: $\arg\min_{\mathbf{S}}\|\Delta\mathbf{Y}\|_{F}^{2}$. Wang et al. [2024] demonstrate that this optimization problem has analytic expression by setting the $\mathbf{S}$ to a lower triangular matrix $\mathbf{L}$, where $\mathbf{L}$ is the Cholesky decomposition of $\mathbf{X}\mathbf{X}^{T}$:
$\mathbf{S}:=\mathbf{L},~{}~{}\text{where}~{}~{}\mathbf{L}\mathbf{L}^{T}=\mathbf{X}\mathbf{X}^{T}.$ (9)
This method takes an additional step to execute the Cholesky decomposition [Meyer, 2000]. Despite this extra computation, it results in a lower output error $\Delta\mathbf{Y}$.
By designing an invertible transformation matrix $\mathbf{S}$, we can transform the weight matrix $\mathbf{W}$ into a decomposition-friendly matrix $\mathbf{WS}$. This transformation takes into account both input and output activations, making the subsequent decomposition more effective for compression. This is so-called Activation-aware Singular Value Decomposition (ASVD).
3.4 Sensitivity-based Truncation Rank Searching
The second challenge arises from the fact that different layers in LLMs exhibit varying degrees of sensitivity to information compression, which is reflected in the distribution of their singular values. Targeting this challenge, we propose the Sensitivity-based Truncation Rank Searching (STRS) method. STRS evaluates the layer sensitivity and decides the best truncation of singular values. In the realm of NLP, perplexity is a key metric for assessing how effectively a language model predicts a sequence of tokens [Brown et al., 2020]. Therefore, we use the reduction in perplexity on the calibration dataset to evaluate the sensitivity of each layer. Similar to post-training compression methods [Dettmers et al., 2022, Frantar et al., 2022, Frantar & Alistarh, 2023], we collect a small number of input token sequences as calibration dataset.
3.5 ASVD for KV Cache Compression